# Active Sparse Bayesian Committee Machine Potential for Isothermal-Isobaric Molecular Dynamics Simulations

### v2

@@ V Conclusions
In this study, we introduced two theoretical advancements to enhance the precision of pressure predictions and to reduce the computational scaling of SGPR MLPs. For the isobaric MD simulations (i.e. $NpT$ and $NpH$), accurate predictions of system pressure is crucial. We showed that the introduction of a virial kernel into the covariance matrix substantially improves pressure predictions, particularly for liquid systems under extreme conditions. This improvement is important for the fidelity of isobaric MD simulations, particularly for the study of phase transition and the behavior under extreme conditions. Furthermore, the refined MLP with the virial kernel enables the estimation of the melting temperature of ice polymorphs. We achieve this by employing $NpH$ MD simulations of the ice-water coexisting phases for three cases (ice II, ice III, and ice V).
@@ Acknowledgements
We are grateful to V. Kapil, S. Pourasad, and C. Schran for helpful discussions. SYW and CWM acknowledge the support from the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00222245). CWM acknowledges the support from the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1C1C1010605). The authors are grateful for the computational support from the Korea Institute of Science and Technology Information (KISTI) for the Nurion cluster (KSC-2021-CRE-0542, KSC-2022-CRE-0115, KSC-2022-CRE-0424, KSC-2023-CRE-0059, KSC-2023-CRE-0198, KSC-2023-CRE-0332, KSC-2023-CRE-0355, KSC-2023-CRE-0454).

### v3

@@ V Conclusions
In this study, we introduced two theoretical advancements to enhance the precision of pressure predictions and to reduce the computational scaling of SGPR MLPs. For the isobaric MD simulations (i.e. $NpT$ and $NpH$), accurate predictions of system pressure is crucial. We showed that the introduction of a virial kernel into the covariance matrix substantially improves pressure predictions, particularly for systems under extreme conditions. This improvement is important for the fidelity of isobaric MD simulations, particularly for the study of phase transition and the behavior under extreme conditions. Furthermore, the refined MLP with the virial kernel enables the estimation of the melting temperature of ice polymorphs. We achieve this by employing $NpH$ MD simulations of the ice-water coexisting phases for three cases (ice II, ice III, and ice V). Additionally, we performed active learning $NpT$ MD simulations of LGPS, demonstrating a phase transition. These results highlight the capability of our improved MLP to accurately capture phase changes in complex materials.
@@ Acknowledgements
We are grateful to V. Kapil, S. Pourasad, and C. Schran for helpful discussions. SYW and CWM acknowledge the support from Brain Pool program funded by the Ministry of Science and ICT through the National Research Foundation of Korea (No. RS-2023-00222245). CWM acknowledges the support from the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1C1C1010605). The authors are grateful for the computational support from the Korea Institute of Science and Technology Information (KISTI) for the Nurion cluster (KSC-2021-CRE-0542, KSC-2022-CRE-0115, KSC-2022-CRE-0424, KSC-2023-CRE-0059, KSC-2023-CRE-0198, KSC-2023-CRE-0332, KSC-2023-CRE-0355, KSC-2023-CRE-0454).